# Through the Judge’s Eyes: Inferred Thinking Traces Improve Reliability of LLM Raters

### v1

@@ 4.3. Fine-Tuning Specialized LLM Rater
We use DeepSeek-R1-0528-Qwen3-8B as our base model (Guo et al., 2025). The model is then fine-tuned on each individual task separately. During fine-tuning, the model is trained not only to predict the final rating but also to generate the entire thought process. This approach encourages the model to learn the underlying reasoning process that leads to a specific judgment. To prevent overfitting, we use early stopping with a patience of 100 steps on a held-out validation set, selecting the model checkpoint that achieves the highest Kendall’s $\tau$ correlation with human ratings. The performance of this final model is then evaluated by comparing its ratings against the human gold standard on the test set.
@@
The results in Table 1 show that reasoning-enhanced SFT significantly improves the performance of the LLM rater. Across all tasks, the models fine-tuned on inferred traces from both DeepSeek-R1 and gpt-oss-120b achieve significant gains over the original model. On average, reasoning-enhanced SFT improves Kendall’s $\tau$ from 0.197 to 0.281, a relative increase of 42.6%. Notably, the final performance of models trained on traces from DeepSeek-R1 and gpt-oss-120b is highly comparable. Both sets of inferred traces, despite originating from different models, serve as effective training data to improve the base rater. This demonstrates the universality of our framework over heterogeneous RLMs.
@@
Metric ICC3 $\tau$ MSE ICC3 $\tau$ MSE ICC3 $\tau$ MSE
Short Story (Complexity) 0.497 0.428 1.116 0.571 0.512 0.634* 0.504 0.401 0.999
Short Story (Engagement) 0.514 0.303 2.044 0.559 0.380 1.525 0.564 0.391* 1.392
Student-Written Essay 0.413 0.421 0.945 0.481 0.491* 0.895 0.433 0.459* 0.957
News Summary (Consistency) 0.150 0.166 2.861 0.178 0.188 2.591 0.181 0.177 2.668
News Summary (Fluency) 0.182 0.142 3.121 0.197 0.192 2.203 0.248 0.190 2.855
Translation 0.419 0.347 2.392 0.448* 0.368 1.942* 0.452 0.356 1.953*
Chatbot Response (Correctness) 0.451 0.354 1.920 0.485 0.370* 1.724* 0.484 0.378* 1.804
Chatbot Response (Helpfulness) 0.475 0.366 1.801 0.511* 0.390* 1.642* 0.484 0.376 1.585*
Average 0.388 0.316 2.025 0.429 0.361 1.645 0.419 0.341 1.777
Table 2. Effect of Codebook Refinement on Human-LLM Rater Agreement. Each value in the table is an average of the corresponding metrics calculated for each LLM rater.
@@
Table 3. Effect of Codebook Refinement on Rater Agreement among LLM Raters (ICC3).
4.4. Refining Annotation Codebook
@@
Crucially, the refined codebooks also improve inter-rater reliability among different LLM raters. Table 3 shows that the average inter-rater agreement (ICC3) increases from 0.580 to 0.621 with the thinking traces inferred by DeepSeek-R1. The benefits are most pronounced on tasks with relative objective criteria, such as correctness of Chatbot Response and Translation. Conversely, on highly subjective tasks like evaluating the engagement of Short Story, the refined codebooks can sometimes increase disagreement between LLM raters. However, this is not necessarily a negative indicator: the original codebook may have induced high but superficial agreement by promoting a shared systematic bias among the LLM raters. By providing more nuanced guidance, the refined codebook reduces this shared bias, causing the models to better reflect the natural and sometimes divergent distribution of human judgments.
@@
Claude (Before):
…The subject story is unrelated to the prompt. It mostly describes a mundane routine with some disjointed details…Then shifts abruptly to snow and vague dialogue about a “gift” and “danger,” but these are underdeveloped. …The narrative lacks structure, coherence, or meaningful plot development. …Elements that could add depth are too vague. …The story reads like disconnected fragments with no complexity. ¡rating¿1¡/rating¿
Claude (After):
…The story includes some elements: narrator routine, home setting, dialogue, and atmosphere. …But development is weak: abrupt jumps, unexplained characters, disconnected dialogue, no plot or character growth, abrupt ending. …Though it introduces multiple ideas (routine, relationships, mysterious “gift”), they are fragmented and underdeveloped. …Falls into the category of “few elements, underdeveloped, fragmented, disconnected.” ¡rating¿2¡/rating¿
DeepSeek (Before):
…The subject story is different from the prompt, lacks magical elements. …Contains surreal and disjointed aspects (time reversing, “gift,” ambiguous dialogue) that add some complexity. …But overall lacks cohesion and depth, feels fragmented. ¡rating¿3¡/rating¿
DeepSeek (After):
…Identifies elements: narrator, teacher, child; mundane events; home/snow setting; vague themes (“gift,” “danger”). …Notes minimal depth, low coherence, abrupt time shifts, unresolved dialogue. …Structure linear but disjointed. …Introduces some elements but underdeveloped, no advanced techniques. …Story is somewhat simple with disconnected elements. ¡rating¿2¡/rating¿
Figure 5. Example thinking traces from two LLM raters on evaluating the complexity of a short story, before and after codebook refinement.
4.5. Ablation Study
@@
To understand the individual contributions of our two-stage refinement process, we conduct an ablation study using thinking traces from DeepSeek R1 on three diverse tasks: Short Story (Engagement), Chatbot Response (Correctness), and News Summary (Consistency). To isolate the effect of each component, we create two separate versions of each codebook: one with only the task instructions refined and another with only the scoring rubric refined, and then assess Human-LLM rater alignment.
Results in Table 4 show that both instruction refinement and rubric refinement enhance performance, though their relative impact depends on the specific weaknesses of the original codebook. For the Chatbot Response task, instruction refinement yields larger gains, increasing Kendall’s $\tau$ from 0.354 to 0.374, compared to 0.359 from rubric refinement. In contrast, rubric refinement proves more effective for News Summary, raising $\tau$ from 0.166 to 0.185, whereas instruction refinement provides only a marginal increase to 0.168. For the Short Story task, both refinements deliver comparably strong improvements. This analysis reveals that the optimal refinement strategy depends on the original codebook’s deficiencies: rubric refinement is most beneficial when the original lacks a detailed rating scale (as with News Summary), while instruction refinement is critical when it lacks a step-by-step guideline, even with a detailed rubric (as with Chatbot Response). These findings validate our two-stage approach, highlighting its robustness in addressing distinct types of flaws in annotation guidelines.
Metric ICC3 $\tau$ MSE ICC3 $\tau$ MSE ICC3 $\tau$ MSE ICC3 $\tau$ MSE
Short Story (Engagement) 0.514 0.303 2.044 0.551 0.372 1.867 0.537 0.380 1.701 0.559 0.380 1.525
Chatbot Response (Correctness) 0.451 0.354 1.920 0.480 0.374 1.827 0.453 0.359 1.836 0.485 0.370 1.724
News Summary (Consistency) 0.150 0.166 2.861 0.145 0.168 2.405 0.195 0.185 2.602 0.178 0.188 2.591
Table 4. Ablation study on the effect of refinement codebook component.
@@
Metric ICC3 $\tau$ MSE ICC3 $\tau$ MSE ICC3 $\tau$ MSE
Short Story (Engagement) 0.514 0.303 2.044 0.559 0.380 1.525 0.533 0.330 1.743
Chatbot Response (Correctness) 0.451 0.354 1.920 0.485 0.370 1.724 0.484 0.376 2.275
News Summary (Consistency) 0.150 0.166 2.861 0.178 0.188 2.591 0.160 0.172 2.533
Table 5. Ablation Study on using post hoc explanation.
@@ 6. Conclusion
Appendix A Dataset Details
@@
The SummEval dataset is a large-scale benchmark designed to evaluate summarization models(Fabbri et al., 2020). The summaries were generated by 23 summarization models trained on CNN/DailyMail news dataset. Each summary is evaluated by crowd-sourced annotators from Amazon Mechanical Turk and three expert annotators, where two had written academic papers on summarization for conferences, and one had completed a senior thesis on the topic. For our analysis, we focus on the expert annotations and retain only the effective samples, defined as those with a standard deviation below 1.0 across expert ratings. This filtering yields 5645 reliable samples out of 6400 total.

### v2

@@ 4.3 Fine-Tuning Specialized LLM Rater
We use DeepSeek-R1-0528-Qwen3-8B and DeepSeek-R1-Distill-Llama-8B as our base model (Guo et al., 2025). The model is then fine-tuned on each individual task separately. During fine-tuning, the model is trained not only to predict the final rating but also to generate the entire thought process. This approach encourages the model to learn the underlying reasoning process that leads to a specific judgment. To prevent overfitting, we use early stopping with a patience of 100 steps on a held-out validation set, selecting the model checkpoint that achieves the highest Kendall’s $\tau$ correlation with human ratings. The performance of this final model is then evaluated by comparing its ratings against the human gold standard on the test set.
@@
The results in Table 1 show that reasoning-enhanced SFT significantly improves the performance of the LLM rater [^4] Due to space limit, we put the results for DeepSeek-R1-Distill-Llama-8B in Appendix A.. Across all tasks, the models fine-tuned on inferred traces from both DeepSeek-R1 and gpt-oss-120b achieve significant gains over the original model. On average, reasoning-enhanced SFT improves Kendall’s $\tau$ from 0.197 to 0.281, a relative increase of 42.6%. Notably, the final performance of models trained on traces from DeepSeek-R1 and gpt-oss-120b is highly comparable. Both sets of inferred traces, despite originating from different models, serve as effective training data to improve the base rater. This demonstrates the universality of our framework over heterogeneous RLMs.
@@ 4.4 Refining Annotation Codebook
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Short Story (Complexity) 0.428 0.497 1.116 0.512 0.571 0.634* 0.401 0.504 0.999
Short Story (Engagement) 0.303 0.514 2.044 0.380 0.559 1.525 0.391* 0.564 1.392
Student-Written Essay 0.421 0.413 0.945 0.491* 0.481 0.895 0.459* 0.433 0.957
News Summary (Consistency) 0.166 0.150 2.861 0.188 0.178 2.591 0.177 0.181 2.668
News Summary (Fluency) 0.142 0.182 3.121 0.192 0.197 2.203 0.190 0.248 2.855
Translation 0.347 0.419 2.392 0.368 0.448* 1.942* 0.356 0.452 1.953*
Chatbot Response (Correctness) 0.354 0.451 1.920 0.370* 0.485 1.724* 0.378* 0.484 1.804
Chatbot Response (Helpfulness) 0.366 0.475 1.801 0.390* 0.511* 1.642* 0.376 0.484 1.585*
Average 0.316 0.388 2.025 0.361 0.429 1.645 0.341 0.419 1.777
Table 2: Effect of Codebook Refinement on Human-LLM Rater Agreement. Each value in the table is an average of the corresponding metrics calculated for each LLM rater.
@@
Table 3: Effect of Codebook Refinement on Rater Agreement among LLM Raters (ICC3).
Figure 6: Effect of Codebook Refinement on Rater Agreement among LLM Raters (ICC3).
@@
Crucially, the refined codebooks also improve inter-rater reliability among different LLM raters. Table 3 shows that the average inter-rater agreement (ICC3) increases from 0.580 to 0.621 with the thinking traces inferred by DeepSeek-R1. The benefits are most pronounced on tasks with relative objective criteria, such as correctness of Chatbot Response and Translation. Conversely, on highly subjective tasks like evaluating the engagement of Short Story, the refined codebooks can sometimes increase disagreement between LLM raters. We hypothesize that this decrease is not necessarily a negative indicator: the original codebook may have induced high but superficial agreement by promoting a shared systematic bias among the LLM raters. By providing more nuanced guidance, the refined codebook might be reducing this shared bias, potentially leading the models to better reflect the natural and sometimes divergent distribution of human judgments.
Claude (Before):
…The subject story is unrelated to the prompt. It mostly describes a mundane routine with some disjointed details…Then shifts abruptly to snow and vague dialogue about a “gift” and “danger,” but these are underdeveloped. …The narrative lacks structure, coherence, or meaningful plot development. …Elements that could add depth are too vague. …The story reads like disconnected fragments with no complexity. <rating>1</rating>
Claude (After):
…The story includes some elements: narrator routine, home setting, dialogue, and atmosphere. …But development is weak: abrupt jumps, unexplained characters, disconnected dialogue, no plot or character growth, abrupt ending. …Though it introduces multiple ideas (routine, relationships, mysterious “gift”), they are fragmented and underdeveloped. …Falls into the category of “few elements, underdeveloped, fragmented, disconnected.” <rating>2</rating>
DeepSeek (Before):
…The subject story is different from the prompt, lacks magical elements. …Contains surreal and disjointed aspects (time reversing, “gift,” ambiguous dialogue) that add some complexity. …But overall lacks cohesion and depth, feels fragmented. <rating>3</rating>
DeepSeek (After):
…Identifies elements: narrator, teacher, child; mundane events; home/snow setting; vague themes (“gift,” “danger”). …Notes minimal depth, low coherence, abrupt time shifts, unresolved dialogue. …Structure linear but disjointed. …Introduces some elements but underdeveloped, no advanced techniques. …Story is somewhat simple with disconnected elements. <rating>2</rating>
Figure 7: Example thinking traces from two LLM raters on evaluating the complexity of a short story, before and after codebook refinement.
@@
4.5 Ablation Study
Task Original Model SFT w/ traces SFT w/o traces
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Short Story (Complexity) 0.177 0.231 2.016 0.463* 0.604* 0.838* 0.329 0.312 1.478
Short Story (Engagement) 0.275 0.289 1.939 0.048 0.364 1.364 0.166 0.182 3.101
Student-Written Essay 0.202 0.233 1.947 0.322* 0.267 1.149* 0.283* 0.277* 2.036
News Summary (Consistency) 0.141 0.186 2.904 0.276 0.319 3.476 0.193 0.236 2.666
News Summary (Fluency) 0.275 0.378 1.938 0.319 0.402 1.160* 0.167 0.381 1.565
Translation 0.171 0.240 2.778 0.260* 0.348* 2.492 0.217 0.272 2.763
Chatbot Response (Correctness) 0.162 0.288 2.049 0.297* 0.403* 1.815* 0.103 0.140 2.742
Chatbot Response (Helpfulness) 0.176 0.274 1.969 0.264* 0.422* 1.723* 0.072 0.120 2.430
Average 0.197 0.265 2.193 0.281 0.391 1.752 0.191 0.240 2.348
Table 4: Ablation Study on the Effect of Thinking Traces in SFT.
Ablation on Thinking Traces in SFT
To disentangle the contribution of reasoning traces from the general benefits of SFT, we trained the same base model using only the final ratings from all available training data across all tasks. Table 4 shows that the model trained exclusively on ratings suffered a performance drop across multiple tasks compared to the trace-augmented model, resulting in overall performance slightly worse than the baseline. This suggests that the model’s inability to utilize intermediate reasoning steps during inference hinders its evaluation capability. Therefore, these results confirm that the generated thinking traces are an essential component for grounding the model’s final judgments.
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Story 0.303 0.514 2.044 0.372 0.551 1.867 0.380 0.537 1.701 0.380 0.559 1.525
Chatbot 0.354 0.451 1.920 0.374 0.480 1.827 0.359 0.453 1.836 0.370 0.485 1.724
Summary 0.166 0.150 2.861 0.168 0.145 2.405 0.185 0.195 2.602 0.188 0.178 2.591
Table 5: Ablation study on the effect of refinement codebook component. Story: Short Story (Engagement). Chatbot: Chatbot Response (Correctness). Summary: News Summary (Consistency).
@@ 4.5 Ablation Study
To understand the individual contributions of our two-stage refinement process, we conduct an ablation study using thinking traces from DeepSeek R1 on three diverse tasks: Short Story (Engagement), Chatbot Response (Correctness), and News Summary (Consistency). To isolate the effect of each component, we create two separate versions of each codebook: one with only the task instructions refined and another with only the scoring rubric refined, and then assess Human-LLM rater alignment. Results in Table 5 show that both instruction refinement and rubric refinement enhance performance, though their relative impact depends on the specific weaknesses of the original codebook. For the Chatbot Response task, instruction refinement yields larger gains, increasing Kendall’s $\tau$ from 0.354 to 0.374, compared to 0.359 from rubric refinement. In contrast, rubric refinement proves more effective for News Summary, raising $\tau$ from 0.166 to 0.185, whereas instruction refinement provides only a marginal increase to 0.168. For the Short Story task, both refinements deliver comparably strong improvements. This analysis reveals that the optimal refinement strategy depends on the original codebook’s deficiencies: rubric refinement is most beneficial when the original lacks a detailed rating scale (as with News Summary), while instruction refinement is critical when it lacks a step-by-step guideline, even with a detailed rubric (as with Chatbot Response). These findings validate our two-stage approach, highlighting its robustness in addressing distinct types of flaws in annotation guidelines.
@@
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Story 0.303 0.514 2.044 0.380 0.559 1.525 0.330 0.533 1.743
Chatbot 0.354 0.451 1.920 0.370 0.485 1.724 0.376 0.484 2.275
Summary 0.166 0.150 2.861 0.188 0.178 2.591 0.172 0.160 2.533
Table 6: Ablation Study on using post hoc explanation.
@@ 6 Conclusion
Acknowledgments
This work was in part supported by a Propelling Original Data Science (PODS) grant sponsored by Michigan Institute for Data & AI in Society (MIDAS) and Microsoft.
Appendix A Additional Tables
Generalizability across Base Architectures
To demonstrate that our framework is not limited to specific model families, we conducted experiments using DeepSeek-R1-Distill-Llama-8B as the base model. Table 7 reports the Human-LLM agreement metrics before and after applying our reasoning-enhanced SFT. The results show consistent improvements across the majority of tasks, confirming that our method effectively generalizes to Llama-based architectures.
Task Original Model SFT (DeepSeek Trace)
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Short Story (Complexity) 0.324 0.419 1.552 0.315 0.419 1.248
Short Story (Engagement) 0.305 0.312 1.199 0.292 0.336 1.108
Student-Written Essay 0.075 0.087 4.475 0.237* 0.260* 1.553*
News Summary (Consistency) 0.059 0.0418 1.782 0.128 0.146 4.018
News Summary (Fluency) 0.159 0.203 1.792 0.213 0.359 1.203*
Translation 0.174 0.143 2.145 0.236* 0.173 2.028*
Chatbot Response (Correctness) 0.063 0.072 2.319 0.154* 0.200* 2.046*
Chatbot Response (Helpfulness) 0.053 0.070 2.434 0.155* 0.164* 2.228*
Average 0.151 0.168 2.212 0.216 0.257 1.929
Table 7: Effect of Reasoning-Enhanced SFT on Human-LLM Rater Agreement (DeepSeek-R1-DistillLlama-8B). For the metrics, higher is better for ICC3 and Kendall’s $\tau$, while lower is better for MSE. An asterisk (*) denotes a statistically significant improvement over the baseline ($p<0.05$).
Sensitivity to Trace Quantity
We further investigated the robustness of our codebook refinement stage regarding the number of thinking traces used. Table 8 presents an ablation study on the Short Story (Engagement) task, varying the number of traces for instruction optimization ($N_{inst}$) and rubric generation ($N_{rubric}$). We observe that performance improves as more traces are utilized, plateauing around our selected configuration ($10/50$). Crucially, even minimal trace settings ($5/20$) outperform the original codebook, indicating that the method is data-efficient and not overly sensitive to hyperparameter tuning.
Setup ($N_{inst}$ / $N_{rubric}$) $\tau$ ICC MSE
Original Codebook 0.303 0.514 2.044
5 / 20 0.370 0.548 1.622
10 / 50 (Selected) 0.380 0.559 1.525
20 / 60 0.396 0.549 1.737
30 / 70 0.363 0.537 1.857
Table 8: Ablation study on the sensitivity of codebook refinement to the number of thinking traces used. We vary the number of traces used for instruction improvement ($N_{inst}$) and rubric improvement ($N_{rubric}$) on the Short Story (Engagement) task. All tested configurations yield better performance than the baseline.
Effectiveness on Generalist Models
To address concerns regarding the necessity of specialized reasoning models, we evaluated our framework on GPT-4o, a strong generalist model without native "thinking" output. Table 9 reports the SFT performance, showing that despite GPT-4o lacking native reasoning traces, our alignment method significantly improves its agreement with human raters (Average $\tau$ improves from 0.197 to 0.237). This trend holds for codebook refinement as well; as shown in Table 10, utilizing traces from GPT-4o yields consistently better evaluation guidelines than the original codebook (Average $\tau$ increasing from 0.316 to 0.339). These results confirm that the benefits of our framework extend to general-purpose LLMs.
Task Original Model SFT (DeepSeek) SFT (gpt-oss) SFT (GPT-4o)
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Story (C) 0.177 0.231 2.016 0.463* 0.604* 0.838* 0.460* 0.581* 0.944* 0.326 0.457* 1.354
Story (E) 0.275 0.289 1.939 0.048 0.364 1.364 0.179 0.371 1.831 0.118 0.421 1.910
Essay 0.202 0.233 1.947 0.322* 0.267 1.149* 0.240 0.244 1.387* 0.259* 0.294* 1.198*
Summary (C) 0.141 0.186 2.904 0.276 0.319 3.476 0.243 0.331 2.970 0.191 0.302 2.388
Summary (F) 0.275 0.378 1.938 0.319 0.402 1.160* 0.306 0.498 1.200* 0.341 0.389 1.627
Translation 0.171 0.240 2.778 0.260* 0.348* 2.492 0.256* 0.314* 2.439* 0.219 0.277 2.404*
Chatbot (C) 0.162 0.288 2.049 0.297* 0.403* 1.815* 0.263* 0.433* 1.897 0.187 0.311 2.085
Chatbot (H) 0.176 0.274 1.969 0.264* 0.422* 1.723* 0.295* 0.466* 1.575* 0.252* 0.430* 1.701*
Average 0.197 0.265 2.193 0.281 0.391 1.752 0.280 0.405 1.780 0.237 0.360 1.833
Table 9: Effect of Reasoning-Enhanced SFT on Human-LLM Rater Agreement (DeepSeek-R1-DistillQwen3-8B) including GPT-4o (non-reasoning LLM).
Task Original Codebook Refined (DeepSeek) Refined (gpt-oss) Refined (GPT-4o)
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Story (C) 0.428 0.497 1.116 0.512 0.571 0.634* 0.401 0.504 0.999 0.387 0.491 1.289
Story (E) 0.303 0.514 2.044 0.380 0.559 1.525 0.391* 0.564 1.392 0.359 0.542 1.956
Essay 0.421 0.413 0.945 0.491* 0.481 0.895 0.459* 0.433 0.957 0.490* 0.492* 0.892
Summary (C) 0.166 0.150 2.861 0.188 0.178 2.591 0.177 0.181 2.668 0.169 0.159 2.904
Summary (F) 0.142 0.182 3.121 0.192 0.197 2.203 0.190 0.248 2.855 0.180 0.169 2.863
Translation 0.347 0.419 2.392 0.368 0.448* 1.942* 0.356 0.452 1.953* 0.362 0.455* 1.882*
Chatbot (C) 0.354 0.451 1.920 0.370* 0.485 1.724* 0.378* 0.484 1.804 0.379 0.481 1.798
Chatbot (H) 0.366 0.475 1.801 0.390* 0.511* 1.642* 0.376 0.484 1.585* 0.385* 0.491 1.530*
Average 0.316 0.388 2.025 0.361 0.429 1.645 0.341 0.419 1.777 0.339 0.410 1.889
Table 10: Effect of Codebook Refinement on Human-LLM Rater Agreement. Each value in the table is an average of the corresponding metrics calculated for each LLM rater.
Appendix B Dataset Details
@@
The SummEval dataset is a large-scale benchmark designed to evaluate summarization models(Fabbri et al., 2021). The summaries were generated by 23 summarization models trained on CNN/DailyMail news dataset. Each summary is evaluated by crowd-sourced annotators from Amazon Mechanical Turk and three expert annotators, where two had written academic papers on summarization for conferences, and one had completed a senior thesis on the topic. For our analysis, we focus on the expert annotations and retain only the effective samples, defined as those with a standard deviation below 1.0 across expert ratings. This filtering yields 5645 reliable samples out of 6400 total.